OpenStudy (anonymous):

suppose you roll two dice. find the probability of rolling a sum of 9.

OpenStudy (anonymous):

we need to count how many ways you can get a total of 9 on two dice it is not that many, so we can list all of them

OpenStudy (anonymous):

\[(4,5), (5,4), (6,3),(3,6)\] are the possible combinations that get a total of 9

OpenStudy (anonymous):

so it would be 4?

OpenStudy (anonymous):

nope, not quite, that is the numerator of our fraction

OpenStudy (anonymous):

there are 4 ways to get a 9, and \(6\times 6=36\) possible rolls all together

OpenStudy (anonymous):

the probability of getting the 9 is therefore \[\frac{4}{36}\] which you can reduce if you like

OpenStudy (anonymous):

i can see if i can find you a table with all the rolls filled in, then these are pretty easy

OpenStudy (anonymous):

not filled in but you can fill it in yourself just add up to get each entry in the 6 by 6 table

OpenStudy (anonymous):

it said the answer was wrong, but it didn't give an explanation as to why.

OpenStudy (anonymous):

maybe it wanted it reduced to \(\frac{1}{9}\) ?

OpenStudy (anonymous):

did you put in 4/36?